# Compositional Scene Understanding through Inverse Generative Modeling

### v2

@@ 2 Related Work
Image Captioning Our work is also related to image captioning. By leveraging pre-trained text-to-image generative models (e.g., Stable Diffusion), our model can be applied to image captioning tasks like BLIP-2 (Li et al., 2023b). However, our approach is applicable to a broader range of scene understanding tasks beyond image captioning. For example, by conditioning on object coordinates, our approach can perform object discovery tasks and even enable generalization to more complex scenes (many more objects) than seen at training. This flexibility and generalizability distinguishes our approach from traditional image captioning models.
@@ 3.1 Compositional Generative Modeling
While this approach is a biased approximation of $p({\bm{x}}|{\bm{c}}^{1},\ldots,{\bm{c}}^{K})$, prior work (Liu et al., 2022) has found that it enables effective compositional generalization to a larger number of visual concepts. We can model each $p({\bm{x}}|{\bm{c}}^{k})$ as an energy-based model (EBM) $e^{-E_{\theta}({\bm{x}}|{\bm{c}}^{k})}$. The product distribution $p({\bm{x}}|{\bm{c}}^{1},\ldots,{\bm{c}}^{K})$ takes the form of a summation of a set of energy functions:
@@
$\epsilon_{\theta}^{\text{comb}}({\bm{x}}^{t},t)=\sum_{k=1}^{K}\epsilon_{\theta}({\bm{x}}^{t},t|{\bm{c}}^{k}),\vspace{-2mm}$ (4)
@@
1:  Input: data distribution $p_{D}$, denoising model $\epsilon_{\theta}$, $\lambda$
@@ 3.2 Compositional Scene Understanding
The inferred set of concepts ${\bm{c}}^{k}$ then corresponds to a description of the scene, where individual concepts can flexibly describe individual objects of the scene as well as global features.
@@ 4.2 Global Factor Perception
Metrics.  We evaluate facial attribute prediction performance with classification accuracy. Classification accuracy is defined as the ratio of correctly classified images to the total number of images, where an image is considered correctly classified only if all the there attributes are simultaneously predicted correctly.
@@ 4.3 Zero-Shot Multi-Object Perception
Diffusion Classifer Variant 67.6%
@@ 5 Limitations and Conclusion
Limitations.  For multiple discrete concept inference, our compositional modeling approach enumerates through each possible configuration for every concept and evaluate denoising error for all concept combinations. This can result in long inference time when the concept number is large. To scale to a large number of concept settings, we developed a continuous approximation of our approach that allows gradient-based optimization in Algorithm 5 and Algorithm 6, thereby avoiding the exponential inference cost. Alternatively, several additional approaches can potentially significantly mitigate this computational bottleneck. One approach could be to use heuristic search algorithms on discrete values – for instance we can run beam search with a beam width of K over each attribute sequentially which can reduce time complexity from $O(M^{K})$ to $O(M*K)$, making inference more efficient for large discrete spaces. Finally, since our approach allows parallel processing across the configurations, inference time can be drastically reduced given sufficient computational resources, potentially approaching the time required for a single configuration evaluation.
@@ Appendix A Appendix
In the Appendix, we first present additional results for various scene understanding tasks in Sec. A.1. Next, we conduct further experiments in Sec. A.2, and provide dataset details in Sec. A.3, training details in Sec. A.4, and baseline details in Sec. A.5. Finally, we discuss potential extensions of our approach in Sec. A.6.
A.1 Additional Results
@@ A.2 Additional Experiments
Prompt Weighting for Multi-Concept Perception.  We conducted an additional comparison to see whether naïve prompt weighting could solve the zero-shot perception task, using the Compel package. Specifically, we applied prompt weighting to the compound prompts including “a photo of a cat++, a dog, and a rabbit”, “a photo of a cat, a dog++, and a rabbit”, and “a photo of a cat, a dog, and a rabbit++”. The underlying idea is that if the image contains specific concepts (e.g., a cat and a dog), then the prompts “a photo of a cat++, a dog, and a rabbit” and “a photo of a cat, a dog++, and a rabbit” are expected to result in lower denoising error (higher likelihood) than the prompt “a photo of a cat, a dog, and a rabbit++”. To determine what two objects present in the scene, we choose the two prompt weighted compound prompts that have the lowest denoising error. We report the zero-shot perception accuracy in Table V. The results suggest that simple prompt weighting may not be sufficient for effective multi-concept inference in this setting.
@@
Object Discovery on CLEVRTex.  Our model has been evaluated on widely adopted datasets (CLEVR and CelebA) that are commonly used for object discovery and multi-label classification. While these datasets are relatively simple, they serve as strong benchmarks for measuring effectiveness and generalization. Nonetheless, we believe that testing our model on more complex scenes would be interesting. To take a first step towards that end, we conducted additional evaluations on ClevrTex (Karazija et al., 2021), which features diverse object colors, textures, shapes, and complex backgrounds. As shown in the Table VI, our model outperforms all baselines in object discovery on ClevrTex, demonstrating its potential scalability to more complex scenarios.
@@
Table VII: Accuracy and Inference Time of Global Factor Perception. The inference time of both our approach and Diffusion Classifer are evaluated on CelebA. To avoid exponential computation cost for discrete concept inference through enumeration, we further develop a continuous approximation of our approach through gradient-based search algorithm, which significant reduces inference time while maintains generalization performance.
@@
Table VIII: Accuracy and Inference Time of Zero-Shot Multi-Object Perception. The inference time of both our approach and Generative Classifer are evaluated on the zero-shot multi-object perception task. To avoid exponential computation cost for discrete concept inference through enumeration, we further develop a continuous approximation of our approach through gradient-based search algorithm, which significantly reduces inference time while maintains generalization performance.
Inference Time.  To evaluate inference efficiency, we conducted a comparison of runtime performance between our method and baseline models on an NVIDIA H100 GPU. The table below shows the inference time for the discrete concept inference on CelebA considering 4 attributes. As shown in Table VII, the inference time of our approach is comparable to the baseline model Diffusion Classifier. To further enable our model to work on a large number of concept settings (i.e., larger K), we developed a continuous approximation of our approach that allows gradient-based optimization, thereby avoiding the exponential ($M^{K}$) inference cost. The gradient-based discrete concept inference algorithm is outlined in Algorithm 5.
Specifically, to infer binary labels with gradient-based optimization, we relax the learnable binary labels to continuous parameters in the range (0, 1). These continuous parameters are optimized using gradient descent and clamped to (0, 1) at each step to remain valid. After optimization, we decide a label is 0 if the corresponding optimized relaxed parameter is smaller than 0.5, otherwise the label is 1. As is shown in the table below, the continuous gradient-based approach reduces inference time significantly, making it scale linearly with the number of concepts ($O(K)$).
@@
A.3 Dataset Details
Object Discovery.  We train our compositional generative model on 50000 CLEVR images (Johnson et al., 2017), each containing 3-5 objects of varying color, shape, size and texture. We resize the training images to a resolution of $64\times 64$. For each image, the 2-D center coordinates of objects are also available, which are of continuous value and normalized between 0 and 1 by dividing the coordinates in terms of pixel value by the image resolution. The out-of-distribution test set consist of images with 6-8 objects either from the CLEVR dataset (Johnson et al., 2017), or from the CLEVRTex dataset (Karazija et al., 2021).

### v4

@@ 2 Related Work
Image Captioning. Our work is also related to image captioning. By leveraging pre-trained text-to-image generative models (e.g., Stable Diffusion), our model can be applied to image captioning tasks like BLIP-2 (Li et al., 2023b). However, our approach is applicable to a broader range of scene understanding tasks beyond image captioning. For example, by conditioning on object coordinates, our approach can perform object discovery tasks and even enable generalization to more complex scenes (many more objects) than seen at training. This flexibility and generalizability distinguishes our approach from traditional image captioning models.
@@ 3.1 Compositional Generative Modeling
While this is a biased approximation, prior work (Liu et al., 2022; Du et al., 2020) has found that it enables effective compositional generalization to a larger number of visual concepts, with a more accurate alternative approximation of $p({\bm{x}}|{\bm{c}}^{1},\ldots,{\bm{c}}^{K})$ discussed in Sec. A.1. We can model each $p({\bm{x}}|{\bm{c}}^{k})$ as an energy-based model (EBM) $e^{-E_{\theta}({\bm{x}}|{\bm{c}}^{k})}$. The product distribution $p({\bm{x}}|{\bm{c}}^{1},\ldots,{\bm{c}}^{K})$ takes the form of a summation of a set of energy functions:
@@
$\epsilon_{\theta}^{\text{comb}}({\bm{x}}^{t},t)=\sum_{k=1}^{K}\epsilon_{\theta}({\bm{x}}^{t},t|{\bm{c}}^{k}),$ (4)
@@
1:  Input: data distribution $p_{D}$, denoising model $\epsilon_{\theta}$
@@ 3.2 Compositional Scene Understanding
The inferred set of concepts ${\bm{c}}^{k}$ then corresponds to a description of the scene, where individual concepts can flexibly describe individual objects as well as global features.
@@ 4.2 Global Factor Perception
Metrics.  We evaluate facial attribute prediction performance with classification accuracy. Classification accuracy is defined as the ratio of correctly classified images to the total number of images, where an image is considered correctly classified only if all the three attributes are simultaneously predicted correctly.
@@ 4.3 Zero-Shot Multi-Object Perception
Diffusion Classifer Variant 73.2%
@@ 5 Limitations and Conclusion
Limitations.  For multiple discrete concept inference, our compositional modeling approach enumerates through each possible configuration for every concept and evaluate denoising error for all concept combinations. This can result in long inference time when the concept number is large. To scale to a large number of concept settings, we developed a continuous approximation of our approach that allows gradient-based optimization in Algorithm 5 and Algorithm 6, thereby avoiding the exponential inference cost. Alternatively, several additional approaches can potentially significantly mitigate this computational bottleneck. One approach could be to use heuristic search algorithms on discrete values – for instance we can run beam search with a beam width of $K$ over each attribute sequentially which can reduce time complexity from $O(M^{K})$ to $O(MK)$, making inference more efficient for large discrete spaces. Finally, since our approach allows parallel processing across the configurations, inference time can be drastically reduced given sufficient computational resources, potentially approaching the time required for a single configuration evaluation.
@@ Appendix A Appendix
In the Appendix, we first present more details regarding the factorization of probability distribution in Sec. A.1. Next, we show additional results for various scene understanding tasks in Sec. A.2, conduct further experiments in Sec. A.3, and provide dataset details in Sec. A.4, training details in Sec. A.5, and baseline details in Sec. A.6. Finally, we discuss potential extensions of our approach in Sec. A.7.
A.1 Distribution Factorization
Assuming conditional independence among concepts ${\bm{c}}^{1},{\bm{c}}^{2},\ldots,{\bm{c}}^{K}$ given ${\bm{x}}$, the distribution $p({\bm{x}}|{\bm{c}}^{1},{\bm{c}}^{2},\ldots,{\bm{c}}^{K})$ can be written as:
$p({\bm{x}}|{\bm{c}}^{1},\ldots,{\bm{c}}^{K})\propto p({\bm{x}},{\bm{c}}^{1},\ldots,{\bm{c}}^{K})=p({\bm{x}})\prod_{k=1}^{K}p({\bm{c}}^{k}|{\bm{x}}).$
In this factorized form, each conditional distribution $p({\bm{c}}^{k}|{\bm{x}})$ can be further approximated by Bayes’ rule as $p({\bm{c}}^{k}|{\bm{x}})\propto\frac{p({\bm{x}}|{\bm{c}}^{k})}{p({\bm{x}})}$, which leads to the following expression:
$p({\bm{x}}|{\bm{c}}^{1},\ldots,{\bm{c}}^{K})\propto p({\bm{x}})\prod_{k=1}^{K}\frac{p({\bm{x}}|{\bm{c}}^{k})}{p({\bm{x}})}.$ (9)
Compared to Eqn (2), the factorization in Eqn (9) involves an additional unconditional term $p({\bm{x}})$. Prior works such as (Liu et al., 2022, 2023) explicitly model this unconditional term, while (Du et al., 2020) assumes that $p({\bm{x}})$ is uniform. In our implementation, we experiment with both approaches and selected the more effective one for each dataset (we include the unconditional term for modeling CLEVR, and use only conditional terms for other datasets).
When incorporating $p({\bm{x}})$ and modeling terms in Eqn (9) as EBMs, the distribution $p({\bm{x}}|{\bm{c}}^{1},\ldots,{\bm{c}}^{K})$ then takes the form:
$p({\bm{x}}|{\bm{c}}^{1},\ldots,{\bm{c}}^{K})\propto e^{-(E_{\theta}({\bm{x}})+\sum_{k=1}^{K}(E_{\theta}({\bm{x}}|{\bm{c}}^{k})-E_{\theta}({\bm{x}})))},$
and the corresponding composed denoising function approximation becomes:
$\epsilon_{\theta}^{\text{comb}}({\bm{x}}^{t},t)=\epsilon_{\theta}({\bm{x}}^{t},t)+\sum_{k=1}^{K}(\epsilon_{\theta}({\bm{x}}^{t},t|{\bm{c}}^{k})-\epsilon_{\theta}({\bm{x}}^{t},t)).$
Based on this approximation, our compositional inverse generative modeling framework can train the compositional diffusion model and estimate likelihood accordingly in the way similar to Eqn (5) and Eqn (7).
A.2 Additional Results
@@ A.3 Additional Experiments
Prompt Weighting for Multi-Concept Perception.  We conducted an additional comparison to see whether naïve prompt weighting could solve the zero-shot perception task, using the Compel package. Specifically, we applied prompt weighting to the compound prompts including “a photo of a cat++, a dog, and a rabbit”, “a photo of a cat, a dog++, and a rabbit”, and “a photo of a cat, a dog, and a rabbit++”. The underlying idea is that if the image contains specific concepts (e.g., a cat and a dog), then the prompts “a photo of a cat++, a dog, and a rabbit” and “a photo of a cat, a dog++, and a rabbit” are expected to result in lower denoising error (higher likelihood) than the prompt “a photo of a cat, a dog, and a rabbit++”. To determine which two objects present in the scene, we choose the two prompt weighted compound prompts that have the lowest denoising error. We report the zero-shot perception accuracy in Table V. The results suggest that simple prompt weighting may not be sufficient for effective multi-concept inference in this setting.
@@
Object Discovery on CLEVRTex.  Our model has been quantitatively evaluated on widely adopted datasets (CLEVR and CelebA) that are commonly used for object discovery and multi-label classification. While these datasets are relatively simple, they serve as strong benchmarks for measuring effectiveness and generalization. Nonetheless, we believe that testing our model on more complex scenes would be interesting. To take a first step towards that end, we conducted additional evaluations on ClevrTex (Karazija et al., 2021), which features diverse object colors, textures, shapes, and complex backgrounds. As shown in the Table VI, our model outperforms all baselines in object discovery on ClevrTex, demonstrating its potential scalability to more complex scenarios.
@@
Table VII: Accuracy and Inference Time of Global Factor Perception. The inference time of both our approach and Generative Classifer are evaluated on CelebA considering 4 attributes including “black hair”, “chubby”, “eyeglasses”, and “smiling”. To avoid exponential computation cost for discrete concept inference through enumeration, we further develop a continuous approximation of our approach through gradient-based search algorithm, which significant reduces inference time while maintains generalization performance.
@@
Table VIII: Accuracy and Inference Time of Zero-Shot Multi-Object Perception. The inference time of both our approach and Diffusion Classifer are evaluated on the zero-shot multi-object perception task. To avoid exponential computation cost for discrete concept inference through enumeration, we further develop a continuous approximation of our approach through gradient-based search algorithm, which significantly reduces inference time while maintains generalization performance.
Inference Time.  To evaluate inference efficiency, we conducted a comparison of runtime performance between our method and baseline models on an NVIDIA H100 GPU. We evaluate inference time for discrete concept inference on CelebA considering more attributes. As shown in Table VII, the inference time of our approach is comparable to the baseline model Generative Classifier. To further enable our model to work on a large number of concept settings (i.e., larger $K$), we developed a continuous approximation of our approach that allows gradient-based optimization, thereby avoiding the exponential ($M^{K}$) inference cost. The gradient-based discrete concept inference algorithm is outlined in Algorithm 5.
Specifically, to infer binary labels with gradient-based optimization, we relax the learnable binary labels to continuous parameters in the range (0, 1). These continuous parameters are optimized using gradient descent and clamped to (0, 1) at each iteration to remain valid. After optimization, we decide a label is 0 if the corresponding optimized relaxed parameter is smaller than 0.5, otherwise the label is 1. As is shown in Table VII, the continuous gradient-based approach reduces inference time significantly, making it scale linearly with the number of concepts ($O(K)$).
@@
A.4 Dataset Details
Object Discovery.  We train our compositional generative model on 26132 CLEVR images (Johnson et al., 2017), each containing 3-5 objects of varying color, shape, size and texture. We resize the training images to a resolution of $64\times 64$. For each image, the 2-D center coordinates of objects are also available, which are of continuous value and normalized between 0 and 1 by dividing the coordinates in terms of pixel value by the image resolution. The out-of-distribution test set consist of images with 6-8 objects either from the CLEVR dataset (Johnson et al., 2017), or from the CLEVRTex dataset (Karazija et al., 2021).